OpenStudy (anonymous):

Two birds start from the same nest and head off in opposite directions. The speed of the first bird is 14 mph more than the speed of the second. After 6 hours the two birds are 348 miles apart. Find the speed of each bird?

OpenStudy (zpupster):

s = d / t speed 1 = x +14 mph speed 2 = x 6 [x + (x+14) ]= 348 6 ( 2x + 14 ) = 348 12x + 84= 348 12x = 348-84 12x = 264 x =22 x+ 14 = 22+14 = 36 Hence, the speed of each bird are 22 mph and 36 mph